# Learning with Incomplete Context: Linear Contextual Bandits with Pretrained Imputation

### v2

@@ 1 INTRODUCTION
Contextual bandits provide a powerful framework for sequential decision-making under uncertainty, where the learner repeatedly observes a context, chooses an action, and receives a reward. The key challenge is to balance exploration and exploitation while adapting decisions to the observed context. Owing to their simplicity and flexibility, contextual bandits have been widely applied in practice, including personalized recommendations (Li et al.,, 2010), mobile health (Nahum-Shani et al.,, 2016), and online education platforms (Cai et al.,, 2021).
In many practical applications, the contexts required for decision-making may be missing or only partially observed during online interactions. For example, in the HeartSteps mobile health study (Liao et al.,, 2020), the full physiological state of a participant is unobserved, while only partial signals such as step counts, activity levels, or self-reports from wearables are available to guide intervention delivery. Similarly, in online education platforms (Lan and Baraniuk,, 2016), a learner’s complete knowledge state across multiple concepts is latent, and the system only observes partial signals such as responses to quiz items or practice problems. At the same time, large offline datasets with substantially more complete contexts are often accessible, since they can be collected without interventions or adaptive decision-making. Such datasets have been shown to reveal richer contextual information than what is available in online interaction (Kausik et al.,, 2025), raising the question of how these auxiliary resources can be effectively leveraged to improve sequential decision-making when online contexts are missing.
In this work, we address the problem of linear contextual bandits when contexts are only partially observed during online interaction, while offline auxiliary data provide full context information. The key idea is to use predictive models trained on auxiliary data to impute the missing contexts for online decisions. Even with access to auxiliary data, it is often reasonable in practice to combine pretrained imputations with simple policies such as linear bandits, since they yield stable and interpretable rules, and enable valid post-hoc statistical inference, which are crucial in applications such as healthcare and education (Rafferty et al.,, 2019; Zhang et al.,, 2024; Guo and Xu,, 2025). Fundamental questions arise: how can predictive models trained on auxiliary data improve such decision rules, and how does imputation quality affect regret?
@@ 1.1 Related works
Bandits with partially observed contexts. Given its importance, a substantial literature has studied contextual bandits with partially observed contexts. Many works impose parametric assumptions on the full context, such as i.i.d. Gaussian contexts or linear dynamical systems with additive Gaussian noise (Kim et al.,, 2023; Park and Faradonbeh,, 2022, 2024; Zeng et al.,, 2025; Xu et al.,, 2021). Others allow more general distributions but with restrictions, such as fixed, time-invariant contexts (Kim et al.,, 2025), or contexts missing completely at random (Jang et al.,, 2022). A closely related work is Hu and Simchi-Levi, (2025), which considers nonlinear bandits with i.i.d. partially observed contexts and leverages pretrained models with orthogonal statistical learning to derive regret bounds. In contrast, we study linear bandits with general contexts that may be dependent, nonstationary, free of parametric assumptions, and missing not at random, and we establish both upper and lower bounds to ensure near-optimality. Another related line of work analyzes corrupted contexts and benchmarks against a mixture of contextual and multi-armed bandits (Bouneffouf,, 2020), whereas our auxiliary data enable comparison to the stronger benchmark of the optimal full-context policy.
Connections to broader areas. Our work also relates to AI-assisted decision-making, where pretrained models support online policies (Tianhui Cai et al.,, 2024; Zhang et al.,, 2025; Chen et al.,, 2021; Janner et al.,, 2021; Lin et al.,, 2023; Lee et al.,, 2023; Ye et al.,, 2025; Cao et al.,, 2024), and to the broad literature on imputation-based methods in statistics and machine learning, from the classical EM algorithm (Dempster et al.,, 1977) to modern ML-based approaches (Xia and Wainwright,, 2024; Angelopoulos et al.,, 2023). We differ by focusing specifically on the missing-context issue in online bandits, providing regret bounds that guide the principled use of pretrained imputation for sequential decisions. A more complete literature review is deferred to the Appendix.
@@ 2 PROBLEM SETUP
    Context generation. A latent context $\bm{Y}_{t}\in\mathbb{R}^{d_{Y}}$ is generated from an unknown probability distribution $p_{\star}(\cdot\mid\bm{Y}_{1:t-1})$. Only a partial observation of $\bm{Y}_{t}$ is revealed to the agent; we denote this observed context by $\bm{S}_{t}\in\mathbb{R}^{d_{S}}$.
@@
Note that in this bandit setting, both $\bm{Y}_{t}$ and its partial observation $\bm{S}_{t}$ are assumed to be exogenous and do not depend on the action sequence $(A_{1},\dots,A_{t})$. This setting naturally arises in many real-world applications. For instance, in digital health interventions, the full state of a patient $\bm{Y}_{t}$ may include physiological and psychological factors such as stress level and sleep quality, while only a subset such as step counts or heart rate ($\bm{S}_{t}$) is observed through wearables and mobile devices. In online education platforms, a learner’s true knowledge state across multiple concepts ($\bm{Y}_{t}$) is unobservable, and the system only receives partial signals like answers to specific quiz items or homework questions.
@@ 3 THE PULSE-UCB ALGORITHM
Under the setting introduced above, we propose Pretrained Unobserved Latent State Estimation UCB (PULSE-UCB), an algorithm that leverages auxiliary data to “fill in the blanks” of the missing contexts before making decisions. The main idea is as follows. We first pretrain a model $\widehat{p}$ on $\mathcal{D}$ that learns to predict the full context $\bm{Y}_{t}$ from the observed sequence $\bm{S}_{1:t}$ [^1] Here $\widehat{p}$ can be any pretrained model that provides a conditional distribution of $\bm{Y}_{t}$ given $\bm{S}_{1:t}$. If $\widehat{p}$ provides a deterministic prediction, one can convert it into a probabilistic model by viewing it as the mean of a suitably chosen distribution. . Then, during online interaction, whenever we only see the partial context $\bm{S}_{1:t}$, we use $\widehat{p}$ to impute the missing parts and obtain complete feature vectors $\widehat{\bm{\phi}}_{t,a}$ for each action. With these surrogate features in hand, the problem reduces to a standard linear contextual bandit, and we apply OFUL (Abbasi-Yadkori et al.,, 2011): the algorithm maintains a confidence set for the unknown parameter $\bm{\theta}^{\star}$, chooses the action that maximizes an optimistic reward estimate, observes the payoff, and updates its estimates accordingly. In this way, the pretrained model provides the missing information, while OFUL handles the exploration-exploitation trade-off.
@@ 4.1 Characterizing imputation error
The next lemma establishes the theoretical basis that a small $D_{t}$ ensures the imputed contexts remain close to the true contexts, enabling reliable downstream decision-making. The proof is in the Appendix.
@@
4.2 Regret analysis under general context
distributions
@@ 4.2 Regret analysis under general context
If $\bm{Y}_{t}$ is conditionally independent of $(\bm{Y}_{1:t-1},\eta_{1:t-1})$ given $\bm{S}_{1:t}$—that is, if $\bm{S}_{1:t}$ provides sufficient information for predicting $\bm{Y}_{t}$—then Assumption 4.1 holds. A simple example is when $\bm{Y}_{t}$ is a function of $\bm{S}_{t}$ with independent randomness. This conditional independence holds naturally in stochastic contextual bandit models (Li et al.,, 2021; Kim et al.,, 2023; Hu and Simchi-Levi,, 2025), where the context at each round $t$ is drawn i.i.d. from a certain distribution. More generally, the assumption also covers broader settings beyond the i.i.d. case.
@@
Ignoring the auxiliary randomness $U_{1:t}$ and considering a simplified setting where $(A_{1:t-1},R_{1:t-1})$, the action and reward prior to round $t$, can be expressed via $(\bm{Y}_{1:t-1},\eta_{1:t-1},\bm{S}_{1:t})$. The term $\mathbb{E}\left[\bm{\Phi}(\bm{Y}_{t},a)\mid\mathcal{G}_{t}\right]$ can then be converted to a conditional expectation over $(\bm{Y}_{1:t-1},\eta_{1:t-1},\bm{S}_{1:t})$. Under Assumption 4.1, such conditional expectation only depends on $\bm{S}_{1:t}$, rendering Equation (4.5) a martingale difference sequence. The complete proof is given in the Appendix.
Remark 4.2.
As a remark, Assumption 4.1 provides a simple and clean framework to handle $\varepsilon_{t}$ with martingale-based tools, enabling regret guarantees. Even without this assumption, however, $\varepsilon_{t}$ can sometimes be controlled by alternative means. For example, if $\bm{W}_{t}$ is generated by a stationary process with geometrically decaying dependence (e.g., a stationary AR process), then concentration inequalities for mixing sequences may be applied for controlling $\varepsilon_{t}$, though such analysis would typically require additional structural assumptions on the feature map $\bm{\Phi}$. Extending Assumption 4.1 to formally cover these dependent settings is left for future work.
With $\varepsilon_{t}$ forming a martingale difference sequence, and considering the reward decomposition (4.4), we can then adapt self-normalized concentration techniques from linear contextual bandits (Abbasi-Yadkori et al.,, 2011). However, an important caveat arises: the effective context $\mathbb{E}\big[\bm{\Phi}(\bm{Y}_{t},A_{t})|\bm{S}_{1:t},A_{t}\big]$ is unknown because the true context distribution $\mathbb{P}$ is unobserved. Instead, it can only be approximated by $\mathbb{E}_{\widehat{p}}\big[\bm{\Phi}(\bm{Y}_{t},A_{t})|\bm{S}_{1:t},A_{t}\big].$ Our analysis therefore requires an additional sensitivity argument that quantifies how inaccuracies in the imputed contexts affect the cumulative regret, leading to regret bounds that explicitly depend on the approximation error between $\widehat{\mathbb{P}}$ and $\mathbb{P}$.
@@
We now state the main result. The next theorem provides a high-probability upper bound on the cumulative regret of Algorithm 1 under general context distributions. The proof is provided in the Appendix.
@@
The proof of both propositions, as well as the explicit choice of $\widehat{p}$ is included in the Appendix. As a remark, the examples above focus on a one-dimensional missing covariate. Extending to a general $d_{W}$-dimensional missing context is straightforward and leads to a similar result, except for an additional $\sqrt{d_{W}}$ factor from handling coordinates separately (e.g., via a union bound or vector-valued concentration).
@@ 5 LOWER BOUNDS
To demonstrate the optimality of our algorithm, we establish a minimax lower bound by analyzing a carefully constructed two-arm contextual bandit instance with partially observed contexts. The exact data-generating process, including feature construction and verification of technical conditions, is provided in the Appendix. We give a high-level overview below.
@@
By alternating between these two regimes, the construction forces both sources of error to matter: historical samples provide noisy information about $f$, while online bandit interaction governs the estimation of $\bm{\theta}^{\star}$. As a result, the minimax regret necessarily includes two additive components—one tied to the nonparametric rate for learning $f$, and the other to the linear rate for estimating $\bm{\theta}^{\star}$. Full details of the construction and proofs are deferred to the Appendix.

### v3

@@ 1 INTRODUCTION
Contextual bandits provide a powerful framework for sequential decision-making under uncertainty, where the learner repeatedly observes a context, chooses an action, and receives a reward. The key challenge is to balance exploration and exploitation while adapting decisions to the observed context. Owing to their simplicity and flexibility, contextual bandits have been widely applied in practice, including personalized recommendations (Li et al., 2010), mobile health (Nahum-Shani et al., 2016), and online education platforms (Cai et al., 2021).
In many practical applications, the contexts required for decision-making may be missing or only partially observed during online interactions. For example, in the HeartSteps mobile health study (Liao et al., 2020), the full physiological state of a participant is unobserved, while only partial signals such as step counts, activity levels, or self-reports from wearables are available to guide intervention delivery. Similarly, in online education platforms (Lan and Baraniuk, 2016), a learner’s complete knowledge state across multiple concepts is latent, and the system only observes partial signals such as responses to quiz items or practice problems. At the same time, large offline datasets with substantially more complete contexts are often accessible, since they can be collected without interventions or adaptive decision-making. Such datasets have been shown to reveal richer contextual information than what is available in online interaction (Kausik et al., 2025), raising the question of how these auxiliary resources can be effectively leveraged to improve sequential decision-making when online contexts are missing.
In this work, we address the problem of linear contextual bandits when contexts are only partially observed during online interaction, while offline auxiliary data provide full context information. The key idea is to use predictive models trained on auxiliary data to impute the missing contexts for online decisions. Even with access to auxiliary data, it is often reasonable in practice to combine pretrained imputations with simple policies such as linear bandits, since they yield stable and interpretable rules, and enable valid post-hoc statistical inference, which are crucial in applications such as healthcare and education (Rafferty et al., 2019; Zhang et al., 2024; Guo and Xu, 2025). Fundamental questions arise: how can predictive models trained on auxiliary data improve such decision rules, and how does imputation quality affect regret?
@@ 1.1 Related works
Bandits with partially observed contexts. Given its importance, a substantial literature has studied contextual bandits with partially observed contexts. Many works impose parametric assumptions on the full context, such as i.i.d. Gaussian contexts or linear dynamical systems with additive Gaussian noise (Kim et al., 2023; Park and Faradonbeh, 2022, 2024; Zeng et al., 2025; Xu et al., 2021). In contrast, we do not impose parametric assumptions on the context distribution, which can be misspecified in complex applications. Other works allow more general distributions but with restrictions, such as fixed, time-invariant contexts (Kim et al., 2025), or contexts missing completely at random (MAR) (Jang et al., 2022). Compared to these works, we allow the context process to be dependent, nonstationary, and missing not at random (MNAR). A closely related work is Hu and Simchi-Levi (2025), which considers nonlinear bandits with i.i.d. partially observed contexts (including both MAR and MNAR scenarios) and leverages pretrained models with orthogonal statistical learning to derive regret upper bounds. Compared to Hu and Simchi-Levi (2025), we focus on linear contextual bandits with potentially dependent and nonstationary contexts and provide both upper and matching lower bounds, establishing near-optimal regret in this setting. Finally, another related line of work analyzes corrupted contexts and benchmarks against a mixture of contextual and multi-armed bandits (Bouneffouf, 2020), whereas our auxiliary data enable comparison to the stronger benchmark of the optimal full-context policy.
Connections to broader areas. Our work also relates to AI-assisted decision-making, where pretrained models support online policies (Tianhui Cai et al., 2024; Zhang et al., 2025; Chen et al., 2021; Janner et al., 2021; Lin et al., 2023; Lee et al., 2023; Ye et al., 2025; Cao et al., 2024), and to the broad literature on imputation-based methods in statistics and machine learning, from the classical EM algorithm (Dempster et al., 1977) to modern ML-based approaches (Xia and Wainwright, 2024; Angelopoulos et al., 2023). We differ by focusing specifically on the missing-context issue in online bandits, providing regret bounds that guide the principled use of pretrained imputation for sequential decisions. A more complete literature review is deferred to the Appendix A.
@@ 2 PROBLEM SETUP
    Context generation. A latent context $\bm{Y}_{t}\in\mathbb{R}^{d_{Y}}$ is generated from an unknown probability distribution $p_{\star}(\cdot\mid\bm{Y}_{1:t-1})$. At each time step $t$, the agent observes a partial context $\bm{S}_{t}\in\mathbb{R}^{d_{S,t}}$. Note that we use the subscript $t$ in the dimension $d_{S,t}$ to explicitly indicate that the dimension of the observed feature is allowed to vary over time.
@@
Note that in this bandit setting, both the full context $\bm{Y}_{t}$ and its partial observation $\bm{S}_{t}$ are assumed to be exogenous and do not depend on the action sequence $(A_{1},\dots,A_{t})$. This asymmetric context availability, where offline auxiliary data provides richer, higher-fidelity information than what can be feasibly observed during real-time online decision-making, is strongly motivated by practical constraints and naturally arises in many real-world applications. For instance, in digital health interventions, the full state of a patient $\bm{Y}_{t}$ may include comprehensive physiological and psychological factors such as stress level and sleep quality. However, because online systems often face strict privacy regulations, local computational limits, or frequent sensor failures, only a partial subset $\bm{S}_{t}$ (such as step counts or heart rate) is observed in real-time. Similarly, in online education platforms, a learner’s true knowledge state across multiple concepts ($\bm{Y}_{t}$) is unobservable during real-time interactions, and the system only receives partial signals like answers to specific quiz items or homework questions ($\bm{S}_{t}$). Fortunately, archival offline data in these domains can safely store the comprehensive state variables ($\bm{Y}_{t}$) to aid online learning. We provide a detailed discussion of these practical motivations in Appendix B.
@@ 3 THE PULSE-UCB ALGORITHM
Under the setting introduced above, we propose Pretrained Unobserved Latent State Estimation UCB (PULSE-UCB), an algorithm that leverages auxiliary data to “fill in the blanks” of the missing contexts before making decisions. The main idea is as follows. We first pretrain a model $\widehat{p}$ on $\mathcal{D}$ that learns to predict the full context $\bm{Y}_{t}$ from the observed sequence $\bm{S}_{1:t}$ [^2] Here $\widehat{p}$ can be any pretrained model that provides a conditional distribution of $\bm{Y}_{t}$ given $\bm{S}_{1:t}$. If $\widehat{p}$ provides a deterministic prediction, one can convert it into a probabilistic model by viewing it as the mean of a suitably chosen distribution. . Then, during online interaction, whenever we only see the partial context $\bm{S}_{1:t}$, we use $\widehat{p}$ to impute the missing parts and obtain complete feature vectors $\widehat{\bm{\phi}}_{t,a}$ for each action. With these surrogate features in hand, the problem reduces to a standard linear contextual bandit, and we apply OFUL (Abbasi-Yadkori et al., 2011): the algorithm maintains a confidence set for the unknown parameter $\bm{\theta}^{\star}$, chooses the action that maximizes an optimistic reward estimate, observes the payoff, and updates its estimates accordingly. In this way, the pretrained model provides the missing information, while OFUL handles the exploration-exploitation trade-off.
@@ 4.1 Characterizing imputation error
The next lemma establishes the theoretical basis that a small $D_{t}$ ensures the imputed contexts remain close to the true contexts, enabling reliable downstream decision-making. The proof is in the Appendix E.
@@
Remark 4.1.
The choice of KL divergence in defining $D_{t}$ in (4.1) is for convenience. In fact, the proof bounds the left-hand side of (4.2) via total variation and Pinsker’s inequality. Thus, any $f$-divergence with a Pinsker-type bound (e.g., Hellinger distance) yields an analogous result. If the left-hand side of (4.2) can be controlled directly, no divergence measure is needed at all.
4.2 Regret analysis under general context distributions
@@
Assumption 4.1 is quite general. It holds whenever $\bm{Y}_{t}$ is conditionally independent of $(\bm{Y}_{1:t-1},\eta_{1:t-1})$ given $\bm{S}_{1:t}$, i.e., whenever $\bm{S}_{1:t}$ provides sufficient information for predicting $\bm{Y}_{t}$. In particular, it is satisfied whenever $\bm{Y}_{t}$ is generated as an arbitrary function of $\bm{S}_{1:t}$ together with independent randomness. This already covers a broader class of models than many existing works on stochastic contextual bandits, which typically assume i.i.d. contexts (Li et al., 2021; Kim et al., 2023; Hu and Simchi-Levi, 2025). Informally, the key requirement is that the new randomness injected at time $t$ does not depend on past randomness; whenever this holds, our algorithm and analysis apply.
@@
An exception arises when $\bm{Y}_{t}$ depends on $\bm{S}_{1:t}$ but is driven by autoregressive (AR) rather than independent noise. In such cases, latent variables exhibit temporal dependence that propagates into the rewards, and different techniques are required. Moreover, different forms of AR structure yield fundamentally different learning dynamics: action-driven AR models (Bacchiocchi et al., 2024) admit modified UCB algorithms with $\mathcal{O}(\sqrt{T})$ regret, whereas latent AR models (Trella et al., 2024) require Kalman-filter-type estimators and may incur regret exceeding $\mathcal{O}(\sqrt{T})$. Even without Assumption 4.1, however, $\varepsilon_{t}$ can sometimes be controlled by alternative means: for example, if $\bm{W}_{t}$ is generated by a stationary process with geometrically decaying dependence (e.g., a stationary AR process), then concentration inequalities for mixing sequences may be applied, though such analysis would typically require additional structural assumptions on the feature map $\bm{\Phi}$. These distinctions underscore that extending beyond Assumption 4.1 to AR settings generally requires case-by-case, model-specific treatment; a detailed discussion is provided in Appendix C.
@@
Ignoring the auxiliary randomness $U_{1:t}$ and considering a simplified setting where $(A_{1:t-1},R_{1:t-1})$, the action and reward prior to round $t$, can be expressed via $(\bm{Y}_{1:t-1},\eta_{1:t-1},\bm{S}_{1:t})$. The term $\mathbb{E}\left[\bm{\Phi}(\bm{Y}_{t},a)\mid\mathcal{G}_{t}\right]$ can then be converted to a conditional expectation over $(\bm{Y}_{1:t-1},\eta_{1:t-1},\bm{S}_{1:t})$. Under Assumption 4.1, such conditional expectation only depends on $\bm{S}_{1:t}$, rendering Equation (4.5) a martingale difference sequence. The complete proof is given in the Appendix F.
With $\varepsilon_{t}$ forming a martingale difference sequence, and considering the reward decomposition (4.4), we can then adapt self-normalized concentration techniques from linear contextual bandits (Abbasi-Yadkori et al., 2011). However, an important caveat arises: the effective context $\mathbb{E}\big[\bm{\Phi}(\bm{Y}_{t},A_{t})|\bm{S}_{1:t},A_{t}\big]$ is unknown because the true context distribution $\mathbb{P}$ is unobserved. Instead, it can only be approximated by $\mathbb{E}_{\widehat{p}}\big[\bm{\Phi}(\bm{Y}_{t},A_{t})|\bm{S}_{1:t},A_{t}\big].$ Our analysis therefore requires an additional sensitivity argument that quantifies how inaccuracies in the imputed contexts affect the cumulative regret, leading to regret bounds that explicitly depend on the approximation error between $\widehat{\mathbb{P}}$ and $\mathbb{P}$.
@@
Remark 4.3 (Alternative Regret Formulation).
Our regret is defined against an oracle observing the full context $Y_{t}$. Alternatively, one might consider a baseline that acts on the conditional expected context: $A_{t}^{\dagger}=\arg\max_{a}\mathbb{E}[\bm{\theta}^{\star\top}\bm{\Phi}(\bm{Y}_{t},a)\mid\bm{S}_{1:t}]$. By Jensen’s inequality, the expected reward of $A_{t}^{\dagger}$ is bounded by that of our full-context oracle:
$\max_{a}\mathbb{E}[\bm{\theta}^{\star\top}\bm{\Phi}(\bm{Y}_{t},a)|\bm{S}_{1:t}]\leq\mathbb{E}\Big[\max_{a}\bm{\theta}^{\star\top}\bm{\Phi}(\bm{Y}_{t},a)|\bm{S}_{1:t}\Big].$
Therefore, bounding the regret against the full-context oracle establishes a strictly stronger theoretical guarantee that naturally upper-bounds the regret against the conditional-expectation baseline.
We now state the main result. The next theorem provides a high-probability upper bound on the cumulative regret of Algorithm 1 under general context distributions. The proof is provided in the Appendix G.
@@
Remark 4.5.
We also provide a variant of Algorithm 1, which combines the imputations with an alternative online exploration algorithm similar to the idea from He et al. (2022). We show that an improved regret guarantee is possible when the order of total imputation error is known and the error varies across time. See Appendix I for details.
Remark 4.6 (Time-Varying Context Dimensions and Structures).
We explicitly emphasize that our framework does not restrict the dimension of the observed state $\bm{S}_{t}$ to be fixed across all time steps. The notation $\bm{S}_{t}\in\mathbb{R}^{d_{S,t}}$ accommodates highly dynamic real-world environments where the availability of sensors or recorded features fluctuates over time (e.g., Missing-Not-At-Random scenarios). Our theoretical guarantees (Theorem 4.1) hold even when the online learner does not share the exact same set of observed features at every round. To empirically substantiate this, we demonstrate PULSE-UCB’s robustness to dynamically changing observed feature sets via a Random Feature Masking experiment in Appendix M.3.
@@
The proof of both propositions, as well as the explicit choice of $\widehat{p}$ is included in the Appendix H. As a remark, the examples above focus on a one-dimensional missing covariate. Extending to a general $d_{W}$-dimensional missing context is straightforward and leads to a similar result, except for an additional $\sqrt{d_{W}}$ factor from handling coordinates separately (e.g., via a union bound or vector-valued concentration). Furthermore, the rates in Propositions 4.1 and 4.2 are robust to mild covariate shift between the historical and online distributions; see Appendix H.3 for details.
@@ 5 LOWER BOUNDS
To demonstrate the optimality of our algorithm, we establish a minimax lower bound by analyzing a carefully constructed two-arm contextual bandit instance with partially observed contexts. The exact data-generating process, including feature construction and verification of technical conditions, is provided in the Appendix J&K. We give a high-level overview below.
@@
By alternating between these two regimes, the construction forces both sources of error to matter: historical samples provide noisy information about $f$, while online bandit interaction governs the estimation of $\bm{\theta}^{\star}$. As a result, the minimax regret necessarily includes two additive components: one tied to the nonparametric rate for learning $f$, and the other to the linear rate for estimating $\bm{\theta}^{\star}$. Full details are deferred to the Appendix J & K.